Update for the Week Ending July 8, 2016

Site Remediation Continues

On-going activities:

  • Continued the application of water and calcium chloride for dust control;
  • Continued the application of odor-suppressant foam and BioSolve for odor and dust control;
  • Maintained erosion and sediment (E&S) controls (Installed waddles - wood chip and straw filled fabric tubes - around the Site perimeter to prevent silt and sediment from leaving the Site and covered all stockpiles);
  • Continued operation of the water treatment plant (WTP) and discharge to the sewer;
  • Continued lining and loading trucks and shipping impacted soil to Bayshore Soil Management for treatment and disposal:
    • 9 trucks; 233.79 tons shipped last week; and
    • 7,606.14 tons shipped to date.
  • Continued installation of the sheet piles in the cement/bentonite (CB) slurry trench for the support of excavation (SOE) wall:
    • 27.6 linear feet installed last week; and
    • 629 of 629 linear feet installed.
  • Initiated installation of dewatering wells inside the SOE wall in preparation for the deep excavation.

Recently completed activities:

  • Completed excavation and installation of the CB slurry trench for installation of the SOE wall sheet piles; and
  • Obtained permitted access to the Bayside property along North 12th Street from NYC Parks and installed temporary fencing for equipment and supplies staging and storage.

Upcoming activities:

  • Complete installation of the sheet piles to the required depth for the SOE wall;
  • Relocate and expand the temporary containment building (TCB) over the SOE wall area in preparation for conducting the deep excavation;
  • Initiate abandonment of monitoring wells located within the excavation; and
  • Initiate the deep excavation within the SOE wall.

Community Air Monitoring Plan Summary

Community air monitoring, initiated on Tuesday, December 1, 2015, continues as intrusive (ground disturbance) activities are conducted.

There were no fifteen-minute Action Limit exceedances for Total Volatile Organic Compounds (TVOCs), Particulate Matter (PM-10) [dust], or odors during this reporting period.

If you have any questions or comments about the work on the Site, please use the Williamsburg MGP Telephone Hotline at (718) 403-3053.

Photos of Activity


Installing sheet piles in the CB slurry trench along the north alignment of the SOE wall.


Installing dewatering wells inside the SOE wall and deep excavation area.


Installing dewatering wells and applying foam to the surrounding area and cuttings for odor control.


Breaking concrete foundations and applying water for dust control.


Installing temporary fencing on the North 12th Street Bayside property for equipment and supplies staging and storage.